# An Infinite Transitivity Theorem

### v2

@@ 1.2. Outline
Finally, we find an irreducible representation of the Calkin algebra which fails infinite transitivity, granting theorem 1. Compare this situation with that of [FA10], wherein the question of whether the commutant of $\mathscr{B}(H)$ in $\mathscr{B}(H)^{\mathscr{U}}$ is trivial also depends on the ultrafilter $\mathscr{U}$. In our case, for a diagonal pure state on $\mathscr{Q}(H)$ (or equivalently $\mathscr{B}(H)$), whether the GNS-representation admits infinite transitivity depends on the corresponding ultrafilter $\mathscr{U}$. Another recent instance of this ultrafilter-dependence phenomenon was in [ZA25].
@@ 2.1. Quantum Filters
There is also an abstract definition of a quantum filter, which is sometimes called a norm filter.
A quantum filter in $A$ is a subset $\mathscr{F}\subseteq A_{+,1}$ with the property that for any $a_{1},\dots,a_{n}\in\mathscr{F}$, $\|a_{1}\cdots a_{n}\|=1$. A maximal quantum filter is a quantum filter not properly contained in any other.
Caution: Quantum filters are not always filters in the order-theoretic sense. More precisely, they are rarely downwards directed under $\ll$ or $\leq$, though interestingly they are always very close [FA19, 5.3.4]. Thankfully, the quantum filters studied in this note will be true (P-)filters under $\ll$.
The beauty of maximal quantum filters is their natural bijective correspondence with pure states, i.e. $\varphi\mapsto\mathscr{F}_{\varphi}$ is a bijection from pure states to maximal quantum filters [FA19, 5.3.8]. However, the map $\varphi\mapsto\mathscr{F}_{\varphi}$ from states to quantum filters is almost never injective or surjective. For example, $\{1\}$ is always a quantum filter, but is only of the form $\mathscr{F}_{\varphi}$ when $A=\mathbb{C}$.
@@ 2.3. Equivalences
Let $\varphi$ be a pure state s.t. $(C,\varphi)$ countably degree-$1$ saturated. Then $\varphi$ has P-excision. If in addition, $C$ has real rank zero, then the excising element can always be chosen to be a projection.
@@
$\|aba-\varphi(b)a^{2}\|\leq\|a\|^{2}\inf_{c\in\mathscr{F}_{\varphi}}\|cbc-\varphi(b)c^{2}\|=0.$
@@
$t(\bar{x})=\{\rho_{j}(P_{j}(\bar{x}))=z_{j}:j\in\mathbb{N}\}.$
Without loss of generality, by the linearity of $\varphi$, whenever $\rho_{j}=\varphi$, assume $z_{j}=0$.
Now fix, for each $m\in\mathbb{N}$, $\bar{b}_{m}$ s.t. $\max_{j\leq m}|\rho_{j}(P_{j}(\bar{b}_{m}))-z_{j}|<\frac{1}{m}.$ Now let $X$ be the collection of entries of each $\bar{b}_{m}$ and coefficients of each $P_{j}$. Then $X$ is countable, so $B=C^{*}(X)$ is separable. Mimicking the proof of 2.9, there exists a single $a\in\mathscr{F}_{\varphi}$ which excises $\varphi$ on $B$. Now define the modified predicate
$\rho^{\prime}_{j}=\begin{cases}\|\cdot\|,&\rho_{j}=\|\cdot\|;\\ \|a\cdot a\|,&\rho_{j}=\varphi;\end{cases}$
and fix the degree-$1$ norm type
$t^{\prime}(\bar{x})=\{\rho^{\prime}_{j}(P_{j}(\bar{x}))=z_{j}:j\in\mathbb{N}\}$
@@ 3. Infinite Transitivity
Let $t(\bar{x})=\{\rho_{j}(P_{j}(\bar{x}))\in K_{j}:j\in\mathbb{N}\}$ be a satisfiable countable type consisting of conditions above. To prove this, we will perform two conversions of $t(\bar{x})$.
Since $t(\bar{x})$ is satisfiable, fix $\bar{b}_{m}$ s.t. $\max_{j\leq m}d(\rho_{j}(P_{j}(\bar{b}_{m})),K_{j})\leq\frac{1}{m}$. Then fix an ultrafilter $\mathscr{U}$ on $\mathbb{N}$. Let $z_{j}=\lim_{m\rightarrow\mathscr{U}}\rho_{j}(P_{j}(\bar{b}_{m}))$. Define
$\sim_{j}=\begin{cases}\leq,&\rho_{j}=\|\cdot\|_{2}^{\varphi};\\ =,&\text{otherwise}.\end{cases}$
Then
$t^{\prime}(\bar{x})=\{\rho_{j}(P_{j}(\bar{x}))\sim_{j}z_{j}:j\in\mathbb{N}\}$
is satisfiable type whose realizations realize $t(\bar{x})$. Now for each $j$ s.t. $\rho_{j}=\varphi$ or $\|\cdot\|_{2}^{\varphi}$, choose $a_{\varphi}\in\mathscr{F}_{\varphi}$ excising the $C^{*}$-subalgebra $B$ generated by coefficients of $P_{j}(\bar{x})$ and entries of $\bar{b}_{m}$. Then, for each $j$ s.t. $\rho_{j}=\varphi$, choose $d_{\varphi}\in B$ s.t. $\varphi(d_{\varphi})=z_{j}$. Define
$P^{\prime}_{j}(\bar{x})=\begin{cases}P_{j}(\bar{x})-d_{\varphi},&\rho_{j}=\varphi;\\ P_{j}(\bar{x}),&\text{otherwise}.\end{cases}$
$\rho^{\prime}_{j}=\begin{cases}\|a_{\varphi}\cdot a_{\varphi}\|,&\rho_{j}=\varphi;\\ \|\cdot a_{\varphi}\|,&\rho_{j}=\|\cdot\|^{\varphi}_{2};\\ \rho_{j},&\text{otherwise}.\end{cases}$
$r_{j}=\begin{cases}0,&\rho_{j}=\varphi;\\ z_{j},&\text{otherwise}.\end{cases}$
@@
$t^{\prime\prime}(\bar{x})=\{\rho^{\prime}_{j}(P^{\prime}_{j}(\bar{x}))\sim_{j}r_{j}:j\in\mathbb{N}\}$
is a satisfiable generalized countable degree-1 norm type whose realizations realize $t^{\prime}(\bar{x})$, thus realize $t(\bar{x})$. To see this, we only need to check the case where $\rho_{j}=\|\cdot\|_{2}^{\varphi}$,
$\|P_{j}(\bar{b})\|_{2}^{\varphi}=\varphi(P_{j}(\bar{b})^{*}P_{j}(\bar{b}))^{\frac{1}{2}}\leq\|a_{\varphi}^{*}P_{j}(\bar{b})^{*}P_{j}(\bar{b})a_{\varphi}\|^{\frac{1}{2}}=\|P_{j}(\bar{b})a_{\varphi}\|\leq r_{j}.$
Therefore, by [FA19, 15.2.10], $t(\bar{x})$ is realized. ∎
We now prove the main results of the paper, infinite Glimm-Kadison for P-representations on coronas of countably degree-1 saturated $C^{*}$-algebras, from which infinite transitivity will follow.
@@
Let $K_{j}$ be a separable subspace of $H_{j}$ and enumerate an orthonormal basis $(\xi_{j})$ of $K=\bigoplus_{j}K_{j}$ compatible with each $K_{j}$. Every $T\in\bigoplus_{j}\mathscr{B}(K_{j})$ is determined by $(T\xi_{j})$. For each $j$, there exists some P-state $\varphi_{j}$ such that there is a lift $a_{j}+L_{\varphi_{j}}=\xi_{j}$. Similarly, there is some lift $b_{j}+L_{\varphi_{j}}=T\xi_{j}$. Define the type
$t(x)=\{\|x\|=\|T\|,\|xa_{j}-b_{j}\|^{\varphi_{j}}_{2}=0:j\in\mathbb{N}\}.$
Since $t(x)$ consists of conditions from 3.1, it suffices to show that it is satisfiable. Fix $F\Subset\mathbb{N}.$ Because $\{\xi_{j}:j\in F\}$ is orthogonal, by the Glimm-Kadison theorem, there exists $d\in C$ s.t. $\pi(d)\xi_{j}=\eta_{j}$ for all $j\in F.$ That is, $\|da_{j}-b_{j}\|^{\varphi}_{2}=0,$ for all $j\in F.$ Note that, because the vectors are uniformly bounded, all these partial realizations can be taken to be in some fixed sort. Therefore $t(x)$ is satisfiable and is realized by $c\in C$, the promised element.
@@ 4. Existence of P-states
Let $A$ be a nonunital $\sigma$-unital $C^{*}$-algebra and $(e_{n})$ a $\sigma$-unit of projections. For any decreasing sequence $X_{n}\subseteq\mathbb{N}$, $a\leq\pi(p_{X_{n}})$ for all $n$ if and only if there exists a pseudointersection $X\subseteq\mathbb{N}$ of $(X_{n})$ s.t. $a\leq\pi(p_{X})$.
@@
The existence of a P-point is extremely common in models of ZFC. In fact, until recently [CH19], there were very few known models of ZFC which have no P-points. Their existence is implied by the continuum hypothesis, is consistent with various standard forcing axioms, and even $\text{OCA}_{\text{T}}$. For the operator algebraist reader who does not fancy the continuum hypothesis, rest assured that P-states are likely relatively consistent which whichever axioms one fancies most.

### v3

@@ 1.2. Outline
Finally, we find an irreducible representation of the Calkin algebra which fails infinite transitivity, granting theorem 1. Compare this situation with that of [FA10], wherein the question of whether the commutant of $\mathscr{B}(H)$ in $\mathscr{B}(H)^{\mathscr{U}}$ is trivial also depends on the ultrafilter $\mathscr{U}$. In our case, for a diagonal pure state on $\mathscr{Q}(H)$ (or equivalently $\mathscr{B}(H)$), whether the GNS-representation admits infinite transitivity depends on the corresponding ultrafilter $\mathscr{U}$. Another recent (and very successful) instance of this ultrafilter-dependence phenomenon was in [ZA25].
@@ 2.1. Quantum Filters
There is also an abstract definition of a quantum filter, namely a quantum filter in $A$ is a subset $\mathscr{F}\subseteq A_{+,1}$ with the property that for any $a_{1},\dots,a_{n}\in\mathscr{F}$, $\|a_{1}\cdots a_{n}\|=1$. A maximal quantum filter is a quantum filter not properly contained in any other.
Caution: Quantum filters are not always filters in the order-theoretic sense. More precisely, they are rarely downwards directed under $\ll$ or $\leq$, though interestingly they are always very close [FA19, 5.3.4] under both orderings. Thankfully, the quantum filters studied in this note will be true (P-)filters under $\ll$.
The beauty of maximal quantum filters is their natural bijective correspondence with pure states, i.e. $\varphi\mapsto\mathscr{F}_{\varphi}$ is a bijection from pure states to maximal quantum filters [FA19, 5.3.8]. However, the map $\varphi\mapsto\mathscr{F}_{\varphi}$ from states to quantum filters is almost never injective or surjective. For example, $\{1\}$ is always a quantum filter, but is only of the form $\mathscr{F}_{\varphi}$ when $A=\mathbb{C}$. In fact, the more natural correspondence is between faces of the state space and quantum filters, though there are still sometimes quantum filters not corresponding to a face of the state space.
@@ 2.3. Equivalences
Let $\varphi$ be a pure state s.t. $(C,\varphi)$ countably degree-$1$ saturated. Then $\varphi$ has P-excision. If in addition $C$ has real rank zero, then the excising element can always be chosen to be a projection.
@@
$\|aba-\varphi(b)a^{2}\|\leq\|a\|^{2}\inf_{c\in Y}\|cbc-\varphi(b)c^{2}\|=0.$
@@
$t(\bar{x})=\{\rho_{j}(P_{j}(\bar{x}))=s_{j}:j\in\mathbb{N}\}.$
By the linearity of $\varphi$, for $\rho_{j}=\varphi$, we can replace $P_{j}(\bar{x})$ by $P_{j}(\bar{x})-d_{j}$, where $d_{j}\in A$ s.t. $\varphi(d_{j})=s_{j}$. Thus, without loss of generality, when $\rho_{j}=\varphi$, assume $s_{j}=0$. Now fix, for each $m\in\mathbb{N}$, $\bar{b}_{m}$ s.t. $\max_{j\leq m}|\rho_{j}(P_{j}(\bar{b}_{m}))-s_{j}|<\frac{1}{m}.$ Now let $X$ be the collection of entries of each $\bar{b}_{m}$ and coefficients of each $P_{j}$. Then $X$ is countable, so $B=C^{*}(X)$ is separable. Mimicking the proof of 2.9, there exists a single $a\in\mathscr{F}_{\varphi}$ which excises $\varphi$ on $B$. Now define the modified *-polynomials
$P^{\prime}_{j}(\bar{x})=\begin{cases}aP_{j}(\bar{x})a,&\rho_{j}=\varphi;\\ P_{j}(\bar{x}),&\rho_{j}=\|\cdot\|;\par\end{cases}$
and fix the countable degree-$1$ norm type
$t^{\prime}(\bar{x})=\{\|P^{\prime}_{j}(\bar{x})\|=s_{j}:j\in\mathbb{N}\}$
@@ 3. Infinite Transitivity
Let $t(\bar{x})=\{\rho_{j}(P_{j}(\bar{x}))\in K_{j}:j\in\mathbb{N}\}$ be a satisfiable countable type consisting of conditions above. To prove that it is realized, we will perform two conversions of $t(\bar{x})$.
Since $t(\bar{x})$ is satisfiable, fix $\bar{b}_{m}$ s.t. $\max_{j\leq m}d(\rho_{j}(P_{j}(\bar{b}_{m})),K_{j})\leq\frac{1}{m}$. Then fix an ultrafilter $\mathscr{U}$ on $\mathbb{N}$. Let $s_{j}=\lim_{m\rightarrow\mathscr{U}}\rho_{j}(P_{j}(\bar{b}_{m}))$. Then
$t^{\prime}(\bar{x})=\{\rho_{j}(P_{j}(\bar{x}))=s_{j}:j\in\mathbb{N}\}$
is satisfiable type whose realizations realize $t(\bar{x})$. Now for each $\varphi$, choose $a_{\varphi}\in\mathscr{F}_{\varphi}$ excising the $C^{*}$-subalgebra $B$ generated by coefficients of $P_{j}(\bar{x})$ and entries of $\bar{b}_{m}$. Then, for each $j$ s.t. $\rho_{j}=\varphi$, choose $d_{j}\in B$ s.t. $\varphi(d_{j})=s_{j}$. Define
$P^{\prime}_{j}(\bar{x})=\begin{cases}a_{\varphi}(P_{j}(\bar{x})-d_{j})a_{\varphi},&\rho_{j}=\varphi;\\ P_{j}(\bar{x})a_{\varphi},&\rho_{j}=\|\cdot\|_{2}^{\varphi};\\ P_{j}(\bar{x}),&\rho_{j}=\|\cdot\|;\end{cases}$
$r_{j}=\begin{cases}0,&\rho_{j}=\varphi;\\ s_{j},&\text{otherwise}.\end{cases}$
@@
$t^{\prime\prime}(\bar{x})=\{\|P^{\prime}_{j}(\bar{x})\|=r_{j}:j\in\mathbb{N}\}$
is a satisfiable countable degree-1 norm type whose realizations realize $t(\bar{x})$. To see this, we only need to check the case where $\rho_{j}=\|\cdot\|_{2}^{\varphi}$,
$\|P_{j}(\bar{b})\|_{2}^{\varphi}=\varphi(P_{j}(\bar{b})^{*}P_{j}(\bar{b}))^{\frac{1}{2}}\leq\|a_{\varphi}^{*}P_{j}(\bar{b})^{*}P_{j}(\bar{b})a_{\varphi}\|^{\frac{1}{2}}=\|P_{j}(\bar{b})a_{\varphi}\|=r_{j}\in K_{j}.$
In this case, $K_{j}$ is of the form $[0,r]$, so $\|P_{j}(\bar{b})\|^{\varphi}_{2}\in K_{j}$. Therefore, since $t^{\prime\prime}(\bar{x})$ is realized, so is $t(\bar{x})$. ∎
We now prove the main results of the paper, infinite Glimm-Kadison for direct sums of P-representations of countably degree-1 saturated $C^{*}$-algebras, from which infinite transitivity will follow.
@@
Let $K_{j}$ be a separable subspace of $H_{j}$ and enumerate an orthonormal basis $(\xi_{i})$ of $K=\bigoplus_{j}K_{j}$ compatible with each $K_{j}$. Every $T\in\bigoplus_{j}\mathscr{B}(K_{j})$ is determined by $(T\xi_{i})$. For each $i,j\in\mathbb{N}$, lift $a_{ij}+L_{\varphi_{j}}=(\xi_{i})_{j}$ and $b_{ij}+L_{\varphi_{j}}=(T\xi_{i})_{j}$. Define the type
$t(x)=\{\|x\|=\|T\|,\|xa_{ij}-b_{ij}\|^{\varphi_{j}}_{2}=0:i,j\in\mathbb{N}\}.$
Since $t(x)$ consists of conditions from 3.1, it suffices to show that it is satisfiable. Fix $m\in\mathbb{N}$. Because $\{\xi_{j}:j\leq m\}$ is orthogonal, by the Glimm-Kadison theorem, for each $\epsilon>0$ there exists $d\in C$ s.t. $\pi_{j}(d)(\xi_{i})_{j}=(T\xi_{i})_{j}$ for all $i,j\leq m$ and $\|d\|<\|T\|+\epsilon$. That is, $\|da_{ij}-b_{ij}\|^{\varphi}_{2}=0,$ for all $j\leq m$ and $\|d\|<\|T\|+\epsilon$. Therefore $t(x)$ is satisfiable and is realized by $c\in C$, the promised element.
@@ 4. Existence of P-states
Let $A$ be a nonunital $\sigma$-unital $C^{*}$-algebra and $(e_{n})$ a $\sigma$-unit of projections. For any decreasing sequence $X_{n}\subseteq\mathbb{N}$ and $a\in\mathscr{Q}(A)_{+,1}$, $a\leq\pi(p_{X_{n}})$ for all $n$ if and only if there exists a pseudointersection $X\subseteq\mathbb{N}$ of $(X_{n})$ s.t. $a\leq\pi(p_{X})$.
@@
The existence of a P-point is extremely common in models of ZFC. In fact, until recently [CH19], there were very few known models of ZFC which have no P-points. Their existence is implied by the continuum hypothesis, is consistent with various standard forcing axioms, and even $\text{OCA}_{\text{T}}$. For the operator algebraist reader who does not fancy the continuum hypothesis, rest assured that P-states are likely relatively consistent with whichever axioms one fancies most.